# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

### v2

@@ Abstract.
We propose a novel algorithm for calculating the ground-state energy of quantum many-body systems by combining auxiliary-field quantum Monte Carlo (AFQMC) with tensor-train sketching. In AFQMC, having a good trial wavefunction to guide the random walk is crucial for avoiding sign problems. Typically, this trial wavefunction is fixed throughout the simulation. Our proposed method iterates between determining a new trial wavefunction in the form of a tensor train, derived from the current walkers, and using this updated trial wavefunction to anchor the next phase of AFQMC. Numerical results demonstrate that our algorithm is highly accurate for large spin systems, achieving a relative error of $10^{-5}$ in estimating ground-state energies. Additionally, the overlap between our estimated trial wavefunction and the ground-state wavefunction achieves a high-fidelity. We provide a convergence proof, highlighting how an effective trial wavefunction can reduce the variance in the AFQMC energy estimate.
@@ 1. Introduction
The quantum many-body problem appears in a wide range of fields, including condensed matter physics, high energy and nuclear physics, quantum chemistry, and material science. One of the most challenging parts of this problem is that the computational cost grows exponentially with the size of the system. Quantum Monte Carlo (QMC) [1, 2, 3, 4] is a class of algorithms that can efficiently deal with high-dimensional problems by reducing the computational cost to a polynomial scale with the system size. It guarantees that in expectation, one can obtain the exact ground-state energy. However, with a finite sample size, QMC generally suffers from sign problem [5, 6], meaning the variance grows exponentially in time. To fix this issue, the constrained-path auxiliary-field quantum Monte Carlo (cp-AFQMC) [7] and phaseless auxiliary-field quantum Monte Carlo (ph-AFQMC) [8] were developed to avoid the sign problem, and have been successfully applied to both lattice models [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] and realistic materials [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. In cp-AFQMC, the wavefunction is represented as a sum of an ensemble of statistical independent random walkers. The sign problem is avoided by introducing a trial wavefunction to bias the random walkers such that they point in the same direction as the trial wavefunction, with a price of a potential systematic bias. This bias depends on the quality of the trial wavefunction [29]. Ideally, the bias will be removed if the trial wavefunction is exactly the ground-state wavefunction of the system [30]. Based on this result, self-consistent approaches for successively improving the trial wavefunction have been developed [31, 32, 33, 34] to reduce the systematic error.
@@
On the other hand, a different type of approach towards the quantum many-body problem is to directly solve for the wavefunction through approximating it by some low-complexity ansatz. One of the most popular ansatz is the matrix product state (MPS) [35, 36], also known as the tensor train (TT) [37]. A huge advantage of TT representation is that its storage complexity and computational complexity can be made near linear (instead of exponential) in the dimension $d$. However, for complicated systems, the approximation error can be large, which leads to inaccurate energy estimates.
@@
Recently, an algorithm that directly combines AFQMC and TT representation has been proposed by one of the authors [38]. In this method, one projects the random walkers into a TT by a TT-sketching technique [39] at every step. It presents a potentially cheaper strategy to update a TT, using AFQMC to simulate a matrix-vector multiplication. While this method avoids the expensive compression of the Hamiltonian operator, it still suffers from the limitation that sometimes one cannot approximate the ground-state wavefunction as a low-rank TT.
@@
In this paper, we provide improvements to AFQMC and tensor methods by combining the best of both worlds. A TT representation of the wavefunction is estimated periodically using TT-sketching. Then the estimated tensor-train wavefunction is used as the trial wavefunction to “anchor” the next episode of AFQMC. Different from a previously proposed method that uses a fixed TT trial wavefunction [40], we alternate between estimating a new trial wavefunction, and running the next episode of AFQMC. While this can potentially remove the energy bias by successively improving the trial wavefunction, we show that this strategy also significantly reduces the energy variance as well. On the other hand, our approach in estimating the tensor-train wavefunction is made cheap by the fact that it does not have to be highly accurate, but just have to be sufficiently accurate to guide the walkers. This is unlike other tensor-train approaches, including [38], where it is imperative to obtain an accurate approximation to the ground state since the tensor-train is used to obtain to final energy estimate. We also motivate our procedure by theoretically showing that an improved trial wavefunction can significantly lower the variance of AFQMC’s energy.
@@
The rest of this paper is organized as follows. In Section 2, we introduce the procedure of cp-AFQMC for calculating the ground state of spin models, and also some preliminaries for TT and TT-sketching. Our proposed algorithm is then presented in Section 3. In Section 4, we provide a theoretical analysis for the convergence of cp-AFQMC, showing that the procedure can give an accurate energy estimate, despite the fact that the wavefunction itself has a large variance. The results of numerical experiments are shown in Section 5.
@@ 2. Preliminary
Applying $e^{-\Delta\tau H}$ to $\Psi^{(n)}$ amounts to applying a $2^{d}\times 2^{d}$ matrix to a vector of size $2^{d}$, which suffers from the curse-of-dimensionality. To solve this issue, QMC performs $e^{-\Delta\tau H}\Psi^{(n)}$ with a Monte Carlo method. In particular, cp-AFQMC represents the wavefunction as the following:
(3) $\Psi^{(n)}\approx\widehat{\Psi}^{(n)}=\sum_{k}{\phi}_{k}^{(n)},$
where each ${\phi}_{k}^{(n)}$ is a meanfield or separable state, and $e^{-\Delta\tau H}$ is decomposed as
@@
While the idea is straightforward, cp-AFQMC further incorporates importance sampling when sampling $\vec{x}_{k}^{(n)}$ and also restricting the path of the random walk in order to reduce the variance of estimating certain quantities using the random walkers. The starting point is a representation of the walkers of the following form
(5) $\widehat{\Psi}^{(n)}=\sum_{k}w_{k}^{(n)}\frac{\phi_{k}^{(n)}}{O_{\mathrm{tr}}(\phi^{(n)}_{k})},$
@@
With this representation, in cp-AFQMC it is possible to zero out some random walkers when $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}\leq 0$, hence constraining the path of AFQMC.
@@
where $\mathcal{N}_{k}^{(n)}$ is the normalization constant. This new probability distribution function favors those walkers that have large overlap with the trial wavefunction.
A remark here is that if an auxiliary field $\vec{x}$ leads to non-positive overlap with the trial function, so that $O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})=0$, then the probability is set to be $\widetilde{P}_{k}^{(n)}(\vec{x})=0$, meaning that this $\vec{x}$ will not be sampled anymore. In this way, we only sample from those $\vec{x}$ such that the walker at the next step $B(\vec{x})\phi_{k}^{(n)}$ has a positive overlap with the trial function, and thus a positive weight. Therefore, the new walker $B(\vec{x})\phi_{k}^{(n)}$ is guaranteed to have a positive weight, as long as $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for at least one $\vec{x}$.
@@
We summarize the algorithm in Algorithm 1.
@@
  • Periodically compute energy as in Section 2.1.1.
  • Periodically do population control as in Section 2.1.2.
@@
In Appendix A.1, we detail the exact implementations for the transverse-field Ising (TFI) model.
2.1.1. Computing ground-state energy
For estimating the ground-state energy, we use the mixed energy estimator
$E_{\text{mixed}}=\frac{\braket{\Psi_{\mathrm{tr}},H\widehat{\Psi}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\widehat{\Psi}^{(n)}}}=\frac{\sum_{k}w^{(n)}_{k}E_{\text{local}}[\Psi_{\mathrm{tr}},\phi_{k}^{(n)}]}{\sum_{k}w^{(n)}_{k}},$
with $\widehat{\Psi}^{(n)}$ in the form of (5), and $E_{\text{local}}$ represents the local energy of a walker:
2.1.2. Population control
As the number of iterations increases, some walkers may accumulate very large weights while some may have very small weights. Those walkers with small weights contribute little to the wavefunction and energy estimation, thus it is meaningless to keep track of them. To make the algorithm more efficient, we periodically implement population control [41] to adjust the weights without changing the statistical distribution. The detailed implementation is explained in Algorithm 5.
@@ 2.2. Tensor-train sketching
which is a separable wavefunction.
The advantage of TT is that it represents a $d$-dimensional tensor with a storage complexity linear in $d$. Moreover, the computational cost of the inner product of two TTs is also linear in $d$, which is numerically verified in Figure 1.
TT-sketching is a procedure to fit some given data into a low-rank TT. In this paper, the data type we consider is a sum of some random walkers as in (5). Each walker is a separable state. A parallel method for this task is given [38], where a left sketching function $S=S_{1}\circ\cdots\circ S_{d}$ and a right sketching function $T=T_{1}\circ\cdots\circ T_{d}$ are used to compress the data. More specifically, denote
@@
where the coefficients $C_{S,j}(\xi_{j-1},l,\xi_{j})$ and $C_{T,j}(\xi_{j-1},l,\xi_{j})$ are Gaussian random tensors, and $\{b_{l}\}$ is a set of basis in the form
@@
We aim to pick a relative small $\delta$ for the basis in (15), such that $\delta^{2}$ is still non-negligible but higher orders of $\delta$ can be ignored, then $S$ and $T$ can be considered as $1$-cluster basis sketching functions [42, 43, 44].
In addition, instead of (13), we form the matrices $X_{j}$ sequentially by making use of the tensor cores we have already obtained, of which the detailed implementation is summarized in Algorithm 2. We remark that in this algorithm, as we sequentially update the tensor cores, the computation for forming $X_{j}$ can be reused, hence the sketching algorithm has a linear time complexity in terms of the dimensionality $d$, which is also verified in Figure 1.
Figure 1. Computational cost of calculating the inner product of a TT trial wavefunction and all the walkers (upper) and TT-sketching (lower). Left: different number of spins with $2000$ walkers; right: different number of walkers for $16$ spins.
@@
      $X_{j}(\xi_{j-1},\gamma_{j-1})=\sum_{\xi_{j-2}}\sum_{\gamma_{j-2}}\sum_{i_{j-1}}X_{j-1}(\xi_{j-2},\gamma_{j-2})S_{j-1}(\xi_{j-2},x_{j-1},\xi_{j-1})\widetilde{G}_{j-1}(\gamma_{j-2},i_{j-1},\gamma_{j-1}),\ j>2;$
@@
      $X_{d}(\xi_{d-1},\gamma_{d-1})=\sum_{\xi_{d-2}}\sum_{\gamma_{d-2}}\sum_{i_{d-1}}X_{d-1}(\xi_{d-2},\gamma_{d-2})S_{d-1}(\xi_{d-2},x_{d-1},\xi_{d-1})\widetilde{G}_{d-1}(\gamma_{d-2},i_{d-1},\gamma_{d-1});$
@@
    $X_{j}(i_{j-1},\gamma_{j-1})\approx\sum_{\alpha_{j-1}=1}^{r_{j-1}}U_{j-1}(\xi_{j-1},\alpha_{j-1})\Sigma_{j-1}(\alpha_{j-1},\alpha_{j-1})V_{j-1}(\alpha_{j-1},\xi_{j-1});$
@@
    $G_{j}(\alpha_{j-1},i_{j},\alpha_{j})=\sum_{\gamma_{j-1}}\sum_{\gamma_{j}}V_{j-1}(\alpha_{j-1},\xi_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})V_{j}(\alpha_{j},\xi_{j}),2\leq j\leq d-1;$
    $G_{1}(i_{1},\alpha_{1})=\sum_{\gamma_{1}}\widetilde{G}_{1}(i_{1},\alpha_{1})V_{1}(\gamma_{1},\alpha_{1}),G_{d}(\alpha_{d-1},i_{d})=\sum_{\gamma_{d-1}}V_{d-1}(\alpha_{d-1},\xi_{d-1})\widetilde{G}_{d}(\gamma_{d-1},i_{d});$
@@ 3. Proposed Algorithm
The constraint $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}>0$ may result in a systematic bias depending on the quality of the trial function $\Psi_{\mathrm{tr}}$. Even in the case where there is no bias, having a better trial wavefunction can improve the variance. To have a better trial wavefunction, we add a step to the cp-AFQMC procedure, where we periodically update the trial function with TT-sketching using the information from the walkers in cp-AFQMC. The procedure is illustrated in Figure 2.
Formally, we summarize our proposed algorithm with details in Algorithm 3.
Figure 2. Proposed framwork. Square nodes represent random walkers, and circle nodes represent the trial wavefunction in a TT form.
Algorithm 3 cp-AFQMC with re-anchoring
@@
An ensemble of walkers $\widehat{\Psi}_{M_{\text{out}}}=\sum_{k}w_{M_{\text{out}},k}\frac{\phi_{M_{\text{out}},k}}{O_{\mathrm{tr}}(\phi_{M_{\text{out}},k})}$, a wavefunction $\Psi_{\mathrm{tr},M_{\text{out}}}$.
Let $\widehat{\Psi}_{0}=\sum_{k}\Psi_{\mathrm{tr}}$ (i.e. $w_{0,k}=1,\phi_{0,k}=\Psi_{\mathrm{tr}}$ which implies $O_{\mathrm{tr}}(\phi_{0,k})=1$).
@@
  • $\widehat{\Psi}_{n_{\text{out}}+1}$ = cp-AFQMC($H,\widehat{\Psi}_{n_{\text{out}}},\Psi_{\mathrm{tr},n_{\text{out}}},\Delta\tau,M_{\text{in}}$).
  • $\Psi_{\mathrm{tr},n_{\text{out}}+1}$ = TT-sketch($\widehat{\Psi}_{n_{\text{out}}+1}$,$S$,$T$).

### v4

@@ Abstract.
We propose a novel algorithm for calculating the ground-state energy of quantum many-body systems by combining auxiliary-field quantum Monte Carlo (AFQMC) with tensor-train sketching. In AFQMC, a good trial wavefunction to guide the random walk is crucial for improving the sampling efficiency and controlling the sign problem. Our proposed method iterates between determining a new trial wavefunction in the form of a tensor train, derived from the current walkers, and using this updated trial wavefunction to anchor the next phase of AFQMC. Numerical results demonstrate that the algorithm is highly accurate for large spin systems. The overlap between the estimated trial wavefunction and the ground-state wavefunction also achieves high fidelity. We additionally provide a convergence analysis, highlighting how an effective trial wavefunction can reduce the variance in the AFQMC energy estimation. From a complementary perspective, our algorithm also extends the reach of tensor-train methods for studying quantum many-body systems.
@@ 1. Introduction
The quantum many-body problem appears in a wide range of fields, including condensed matter physics, high energy and nuclear physics, quantum chemistry, and material science. One of the most challenging parts of this problem is that the computational cost grows exponentially with the size of the system. Quantum Monte Carlo (QMC) [1, 2, 3, 4] is a class of algorithms that can efficiently deal with high-dimensional problems by reducing the computational cost to a polynomial scale with the system size. It guarantees that, in expectation, one can obtain the exact ground-state energy. However, with a finite sample size, QMC generally suffers from the sign problem [5, 6], meaning the variance grows exponentially in time. To fix this issue, the constrained-path auxiliary-field quantum Monte Carlo (cp-AFQMC) [7] and phaseless auxiliary-field quantum Monte Carlo (ph-AFQMC) [8] were developed to control the sign problem, and have been successfully applied to both lattice models [9, 10, 11, 12, 13, 14, 15, 16, 17] and realistic materials [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In cp-AFQMC, the wavefunction is represented formally as the sum of an ensemble of statistically independent random walkers. The sign problem is avoided by introducing a trial wavefunction to bias the random walkers such that they maintain a non-negative overlap or a constant gauge with the trial wavefunction, with the price of a potential systematic error. This bias depends on the quality of the trial wavefunction [28]. Ideally, the bias will be removed if the trial wavefunction is exactly the ground-state wavefunction of the system [29]. Based on this result, self-consistent approaches for successively improving the trial wavefunction have been developed [30, 31, 32, 33] to reduce the systematic error.
@@
On the other hand, a different type of approach towards the quantum many-body problem is to directly solve for the wavefunction through approximating it by some low-complexity ansatz. One of the most popular ansatz is the matrix product state (MPS) [34, 35], also known as the tensor train (TT) [36]. A huge advantage of TT representation is that its storage complexity and computational complexity can be made near linear (instead of exponential) in the dimension $d$. However, for complicated systems, the approximation error can be large, which leads to inaccurate estimates of physical observables.
@@
Recently, an algorithm that directly combines AFQMC and TT representation has been proposed by one of the authors [37]. In this method, the random walkers are projected into a TT by a TT-sketching technique [38] at every step. It presents a potentially cheaper strategy to update a TT, using AFQMC to simulate a matrix-vector multiplication. Although this method avoids the expensive compression of the Hamiltonian operator, it still suffers from the limitation that sometimes one cannot approximate the ground-state wavefunction as a low-rank TT.
@@
In this paper, we provide improvements to AFQMC and tensor methods by combining the best of both worlds. A TT representation of the wavefunction is estimated periodically using TT-sketching. Then the estimated tensor-train wavefunction is used as the trial wavefunction to “anchor” the next episode of AFQMC. Different from a previously proposed method that uses a fixed TT trial wavefunction [39], our method iteratively builds the TT trial wavefunction from the AFQMC walkers. Specifically, we alternate between estimating a new trial wavefunction, and running the next episode of AFQMC. This can potentially remove the bias in cases where the sign/phase problem is present, by successively improving the trial wavefunction. It can also significantly reduce the variance. Our approach in estimating the tensor-train wavefunction is efficient, as it requires only enough accuracy to effectively guide the walkers. This is unlike other tensor-train approaches, including [37], where it is imperative to obtain an accurate approximation to the ground state since the tensor-train is used to obtain the final energy estimate.
@@
The rest of this paper is organized as follows. In Section 2, we introduce the procedure of cp-AFQMC for calculating the ground state of spin models, and also some preliminaries for TT and TT-sketching. Our proposed algorithm is then presented in Section 3. In Section 4, we provide a theoretical analysis for the convergence of cp-AFQMC, showing that the procedure can give an accurate energy estimate, despite the fact that the wavefunction itself can have large variance. The results of numerical experiments are shown in Section 5.
@@ 2. Preliminary
Applying $e^{-\Delta\tau H}$ to $\Psi^{(n)}$ amounts to applying a $2^{d}\times 2^{d}$ matrix to a vector of size $2^{d}$, which suffers from the curse of dimensionality. To solve this issue, QMC performs $e^{-\Delta\tau H}\Psi^{(n)}$ with a Monte Carlo method. In particular, cp-AFQMC represents the wavefunction as the following:
(3) $\Psi^{(n)}\approx\widehat{\Psi}^{(n)}=\sum_{k=1}^{N}{\phi}_{k}^{(n)},$
where each ${\phi}_{k}^{(n)}$ is a meanfield or separable state, $N$ is the number of walkers, and $e^{-\Delta\tau H}$ is decomposed as
@@
As key algorithmic ingredients, cp-AFQMC further incorporates importance sampling when sampling $\vec{x}_{k}^{(n)}$in order to reduce the variance, which will also constrain the paths of the random walk when the sign problem is present. The starting point is a representation of the walker distribution of the following form
(5) $\widehat{\Psi}^{(n)}=\sum_{k=1}^{N}w_{k}^{(n)}\frac{\phi_{k}^{(n)}}{O_{\mathrm{tr}}(\phi^{(n)}_{k})},$
@@
With this representation, in cp-AFQMC it is possible to zero out some random walkers when $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}\leq 0$, hence constraining the paths.
@@
where $\mathcal{N}_{k}^{(n)}$ is the normalization constant. This new probability distribution function favors those walkers that have a large overlap with the trial wavefunction.
A remark here is that if an auxiliary field $\vec{x}$ leads to non-positive overlap with the trial function, so that $O_{\mathrm{tr}}(B(\vec{x})\phi_{k}^{(n)})=0$, then the probability is set to $\widetilde{P}_{k}^{(n)}(\vec{x})=0$, meaning that this $\vec{x}$ will not be sampled anymore. In this way, we only sample from those $\vec{x}$ such that the walker at the next step $B(\vec{x})\phi_{k}^{(n)}$ has a positive overlap with the trial function, and thus a positive weight. Therefore, the new walker $B(\vec{x})\phi_{k}^{(n)}$ is guaranteed to have a positive weight, as long as $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for at least one $\vec{x}$.
@@
To estimate the ground-state energy, we use the mixed estimator
(13) $E_{\text{mixed}}=\frac{\braket{\Psi_{\mathrm{tr}},H\widehat{\Psi}^{(n)}}}{\braket{\Psi_{\mathrm{tr}},\widehat{\Psi}^{(n)}}}=\frac{\sum_{k}w^{(n)}_{k}E_{\text{local}}[\Psi_{\mathrm{tr}},\phi_{k}^{(n)}]}{\sum_{k}w^{(n)}_{k}},$
where $E_{\text{local}}$ represents the local energy of a walker:
We summarize the whole algorithm in Algorithm 1.
@@
  • Periodically estimate the ground-state energy using the mixed estimator (13).
  • Periodically do population control as in Algorithm 5 in the appendix.
@@
In Appendix C.1, we detail the exact implementations for the transverse-field Ising (TFI) model.
@@ 2.2. Tensor-train sketching
which is a separable wavefunction. The advantage of TT is that it represents a $d$-dimensional tensor with a storage complexity linear in $d$. Moreover, the computational cost of the inner product of two TTs is also linear in $d$, which is numerically verified in Figure 12.
TT-sketching is a procedure to fit some given data into a low-rank TT. In this paper, the data type we consider is a sum of random walkers as in (5). Each walker is a separable state. A parallel method for this task is given [37], where a left sketching function $S=S_{1}\circ\cdots\circ S_{d}$ and a right sketching function $T=T_{1}\circ\cdots\circ T_{d}$ are used to compress the data. More specifically, denote
@@
where the coefficients $C_{S,j}(\xi_{j-1},l,\xi_{j})$ and $C_{T,j}(\xi_{j-1},l,\xi_{j})$ are Gaussian random tensors, and $\{v_{l}\}$ is a set of basis in the form
@@
We aim to pick a relative small $\delta$ for the basis in (16), such that $\delta^{2}$ is still non-negligible but higher orders of $\delta$ can be ignored. Then $S$ and $T$ can be considered as $1$-cluster basis sketching functions [40, 41, 42].
In addition, instead of (14), we form the matrices $X_{j}$ sequentially by making use of the tensor cores we have already obtained, the detailed implementation of which is summarized in Algorithm 2. The tensor diagram illustrating the TT-sketching procedure is also provided in Figure 1. We remark that in this algorithm, as we sequentially update the tensor cores, the computation for forming $X_{j}$ can be reused, so the sketching algorithm has a linear time complexity in terms of the dimensionality $d$, which is verified in Figure 13 in the appendix.
@@
      $X_{j}(\xi_{j-1},\gamma_{j-1})=\sum_{\xi_{j-2}}\sum_{\gamma_{j-2}}\sum_{i_{j-1}}X_{j-1}(\xi_{j-2},\gamma_{j-2})S_{j-1}(\xi_{j-2},i_{j-1},\xi_{j-1})\widetilde{G}_{j-1}(\gamma_{j-2},i_{j-1},\gamma_{j-1}),\ j>2;$
@@
      $X_{d}(\xi_{d-1},\gamma_{d-1})=\sum_{\xi_{d-2}}\sum_{\gamma_{d-2}}\sum_{i_{d-1}}X_{d-1}(\xi_{d-2},\gamma_{d-2})S_{d-1}(\xi_{d-2},i_{d-1},\xi_{d-1})\widetilde{G}_{d-1}(\gamma_{d-2},i_{d-1},\gamma_{d-1});$
@@
    $X_{j}(\xi_{j-1},\gamma_{j-1})\approx\sum_{\alpha_{j-1}=1}^{r_{j-1}}U_{j}(\xi_{j-1},\alpha_{j-1})\Sigma_{j}(\alpha_{j-1},\alpha_{j-1})V_{j}(\alpha_{j-1},\gamma_{j-1});$
@@
    $G_{j}(\alpha_{j-1},i_{j},\alpha_{j})=\sum_{\gamma_{j-1}}\sum_{\gamma_{j}}V_{j}(\alpha_{j-1},\gamma_{j-1})\widetilde{G}_{j}(\gamma_{j-1},i_{j},\gamma_{j})V_{j+1}(\alpha_{j},\gamma_{j}),2\leq j\leq d-1;$
    $G_{1}(i_{1},\alpha_{1})=\sum_{\gamma_{1}}\widetilde{G}_{1}(i_{1},\alpha_{1})V_{2}(\gamma_{1},\alpha_{1}),G_{d}(\alpha_{d-1},i_{d})=\sum_{\gamma_{d-1}}V_{d}(\alpha_{d-1},\gamma_{d-1})\widetilde{G}_{d}(\gamma_{d-1},i_{d});$
Figure 1. Tensor diagrams illustrating the TT-sketching described in Algorithm 2. As a demonstration, we present the steps involved constructing the third tensor core $G_{3}$ for a system with $d=6$. We refer the reader to [43] for a detailed introduction to tensor network diagrams.
@@ 3. Proposed Algorithm
In AFQMC, if there is a sign problem, the constraint $\braket{\Psi_{\mathrm{tr}},\phi_{k}^{(n)}}>0$ may result in a systematic error, which depends on the quality of the trial function $\Psi_{\mathrm{tr}}$. Even in the case where there is no sign problem, a better trial wavefunction can still reduce the variance of the random walk and improve sampling efficiency. To improve the trial wavefunction, our algorithm periodically updates the trial function with TT-sketching using the information from the walkers in cp-AFQMC. The procedure is illustrated in Figure 2. Formally, we summarize our proposed algorithm with details in Algorithm 3.
[…]
Figure 2. Proposed framework: cp-AFQMC-re-anchoring.
Algorithm 3 cp-AFQMC-re-anchoring
@@
An ensemble of walkers $\widehat{\Psi}^{(M_{\text{in}}M_{\text{out}})}=\sum_{k}w^{(M_{\text{in}}M_{\text{out}})}_{k}\frac{\phi^{(M_{\text{in}}M_{\text{out}})}_{k}}{O_{\mathrm{tr}}(\phi^{(M_{\text{in}}M_{\text{out}})}_{k})}$, a wavefunction $\Psi_{\mathrm{tr},M_{\text{out}}}$.
Let $\widehat{\Psi}^{(0)}=\sum_{k}\Psi_{\mathrm{tr,0}}$ (i.e. $w_{0,k}=1,\phi_{0,k}=\Psi_{\mathrm{tr}}$ which implies $O_{\mathrm{tr}}(\phi_{0,k})=1$).
@@
  • $\widehat{\Psi}^{(M_{\text{in}}n_{\text{out}}+M_{\text{in}})}$ = cp-AFQMC($H,\widehat{\Psi}^{(M_{\text{in}}n_{\text{out}})},\Psi_{\mathrm{tr},n_{\text{out}}},\Delta\tau,M_{\text{in}}$).
  • $\Psi_{\mathrm{tr},n_{\text{out}}+1}$ = TT-sketching($\widehat{\Psi}^{(M_{\text{in}}n_{\text{out}}+M_{\text{in}})}$,$S$,$T$).